The average number of milligrams (mg) of sodium in a certain brand of low-salt microwave frozen dinners is 660 mg, and the standard deviation is 35 mg. Assume the variable is normally distributed. If a sample of 10 dinners is selected, find the probability that the mean of the sample will be larger than 670 mg.
Answer by stanbon(75887)

The average number of milligrams (mg) of sodium in a certain brand of low-salt microwave frozen dinners is 660 mg, and the standard deviation is 35 mg. Assume the variable is normally distributed. If a sample of 10 dinners is selected, find the probability that the mean of the sample will be larger than 670 mg.
------------------
t(670) = (670-660)/[35/sqrt(10)] = 0.9035
-------
P(x-bar > 670) = P(t > 0.9035 when df = 9) = tcdf(0.9035,100,9) = 0.1949
-----------------
Cheers,
Stan H.
---------------